OpenStudy (anonymous):

@LagrangeSon678 can you help with my problem

OpenStudy (anonymous):

whats the problem?

OpenStudy (anonymous):

Given the linear equation y=-2/7x+3, find the y-coordinates of the points (-7, ), (0, ), and (7, ). Please show all of your work

OpenStudy (anonymous):

You have y=-2/7x+3, and you are given three x-values,-7,0,and 7. What you should do is plug in each one of these x values into your given equation, like so: y=-2/7(-7)+3 y=-2/7(0)+3 y=-2/7(7)+3 Evaluateing these will produce the correspoding missing y-values

OpenStudy (anonymous):

then what do i do from there that confuses me

OpenStudy (anonymous):

replace what you get in the ordered pairs, (-7, ), (0, ), and (7, )

OpenStudy (anonymous):

For instance i will do one for you, given (-7, ) i will replace the x in your equation with -7, this gives me: y=-2/7(-7)+3 =2+3 =5 thus y=5, now simply we have (-7,5)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

y=2/7(7)+3?

OpenStudy (anonymous):

you mean y=-2/7(7)+3, you missed the negative sign, regardless this will give you y=1, (7,1)

OpenStudy (anonymous):

then for x=0, y=-2/7(0)+3--->y=3 (0,3)

OpenStudy (anonymous):

SO THE NUMBER YOU SUB IS THE NUMBER YOU PUT WIT THE THREE

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

NOW HOW DID YOU GET THE 1 IN THE PREVIOUS PROBLEM

OpenStudy (anonymous):

for (-7, ) ?

OpenStudy (anonymous):

FOR THE 7

OpenStudy (anonymous):

y=-2/7(7)+3 =-2+3 =-1 Thus, (7,-1)

OpenStudy (anonymous):

SO WHAT HAPPENS TO THE TWO 7'S

OpenStudy (anonymous):

you were given three ordered pair, like this (x,y). Excpet you where only given the x value. (-7, ?) (0,?) (7,?) So this means that inorder to find the y value, you have to plug the x values into your equation y=-2/7x+3.

OpenStudy (anonymous):

RIGHT I GET THAT BUT WHAT HAPPENS WHEN YOU YOU PLUGGED IN THE 7 HOW DID YOU GET TO THE -1

OpenStudy (anonymous):

y=-2/7x+3, this is your equation, y=-2/7(7)+3 =-2+3 1

OpenStudy (anonymous):

OKAY I SEE NOW THANKS SO MUCH

OpenStudy (anonymous):

your welcome